OpenStudy (anonymous):

@AnswerMyQuestions @Directrix @iGreen FIND THE SUM OR DIFFRENCE 7/8+9/16 A.2/3 B.1 1/2 C.1 7/16

OpenStudy (anonymous):

What is the smallest number 8 and 16 can go into?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

No...

OpenStudy (anonymous):

You have to multiply.

OpenStudy (igreen):

No, can 8 get into 16?

OpenStudy (anonymous):

YES

OpenStudy (igreen):

Yes, so that will be our LCD.

OpenStudy (igreen):

Multiply 2 to 7/8 2 * 7 = ? 2 * 8 = ?

OpenStudy (anonymous):

14 16

OpenStudy (igreen):

Yep

OpenStudy (igreen):

14 + 9 = ?

OpenStudy (anonymous):

23

OpenStudy (igreen):

Yep, that gives us 23/16. Now we have to change it into a Mixed Number.

OpenStudy (igreen):

 See how I did that?